OpenStudy (anonymous):

what is the integral (x/(x+4) )dx ????/

OpenStudy (zarkon):

\[\frac{x}{x+4}=\frac{x+4-4}{x+4}=\frac{x+4}{x+4}-\frac{4}{x+4}=1-\frac{4}{x+4}\]